OpenStudy (anonymous):

the length of a rectangle is twice the width the area is 98 yd^2 find the length and the width

OpenStudy (anonymous):

If w=width, then the length is 2w. Length (2w) * Width (w) = Area (A), so 2w*w=98. \[2w ^{2} = 98\]. 98/2 =49. \[\sqrt{49}\] = 7. So then the width is 7, and the length (2 * 7) is 14.

OpenStudy (anonymous):

ty